OpenStudy (anonymous):

Jarvis and his family took a road trip. His dad drove for the first half of the trip. His mom drove for 2/3 of the remaining time, and then Jarvis drove the rest of the way. If the total trip took 10 hours, how long did Jarvis drive?

OpenStudy (texaschic101):

if his dad drove for the first half of the trip, and the trip was 10 hrs, then his dad drove 5 hrs. Leaving 5 hrs. And his mom drove for 2/3 of 5 hrs, which is 2/3 * 5 = 10/3 = 3 1/3 hrs. Leaving Jarvis to drive : 5 - 3 1/3 = 5 - 10/3 = 15/3 - 10/3 = 5/3 = 1 2/3 hrs check.. 5 + 10/3 + 5/3 = 15/3 + 10/3 + 5/3 = 30/3 = 10 hrs (correct) Jarvis drove for 1 2/3 hrs